OpenStudy (anonymous):

Write a rule for the linear function in the table. x:1,2,3,4 f(x):3,2,1,0 A. f(x) = 4x B. f(x) = x – 4 C. f(x) = x + 4 D. f(x) = 4 – x

OpenStudy (hoblos):

f(x) = 4 – x 4-1=3 4-2=2 4-3=1 4-4=0